OpenStudy (anonymous):

The polynomial of degree 4, P(x) has a root of multiplicity 2 at x=2 and roots of multiplicity 1 at x=0 and x=−3. Its lead coefficient is 4. Find a formula for P(x).

OpenStudy (johnweldon1993):

multiplicity of 2...means ² multiplicity of 1 means just ^1 so if the roots are at 2, 0, and -3 (x -2)(x-0)(x+3) your leading coefficent goes in front 4(x -2)(x-0)(x+3) now which has which multiplicity? 4(x -2)²(x-0)(x+3) so your P(X) = 4(x -2)²(x-0)(x+3) if you want to simplify...go ahead

OpenStudy (anonymous):

thanks can you help me with this last one please Use the Factor Theorem to find a factor of P(x)=x^3−6x^2+10x−5. as follows: Compute P(1) = and P(−1) = . Conclude that________ is a factor of P(x)=x3−6x2+10x−5 i just need the last part